OpenStudy (anonymous):

How to integrate ((x^2)-3x+7)/((x^2)-4x+6)^2 Please someone help me find the solution..

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{x^2-3x+7}{x^2-4x+6}~dx}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{x^2-4x+x+6+1}{x^2-4x+6}~dx}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{x^2-4x+6}{x^2-4x+6}+\frac{x+1}{x^2-4x+6}~dx}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}1+\frac{x+1}{x^2-4x+6}~dx}\)

OpenStudy (solomonzelman):

This is where I would start, doing a division but that way to make it faster.

OpenStudy (anonymous):

The denominator have the power

OpenStudy (solomonzelman):

oh,

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{x^2-3x+7}{\left(x^2-4x+6\right)^2}~dx}\)

OpenStudy (solomonzelman):

well, if I do my method I don't know if that is helpful or not you would have: \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{1}{\left(x^2-4x+6\right)}+\frac{x+1}{\left(x^2-4x+6\right)^2}dx}\)

OpenStudy (solomonzelman):

I checked the answer it is rediculous, but it is findable.... What have you attempted so far?

OpenStudy (anonymous):

The answer is (7/8)(surd 2)tan inverse((x-2)/surd 2)+((3x-8)/((4x^2)-16x+24)+C

OpenStudy (solomonzelman):

curd?

OpenStudy (solomonzelman):

surd * ?

OpenStudy (solomonzelman):

The wolfram gives this, but that is a long work to get to.... http://www.wolframalpha.com/input/?i=Integral+%28x%5E2-3x%2B7%29%2F%28%28x%5E2-4x%2B6%29%5E2%29 (I assume I am not violating the policy by posting this)

OpenStudy (solomonzelman):

if you start from, then you know integral#1 by completing the square. \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{1}{\left(x^2-4x+6\right)}+\frac{x+1}{\left(x^2-4x+6\right)^2}dx}\) and integral#2 is what we should figure out.

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{x+1}{\left(x^2-4x+6\right)^2}dx}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{x+1}{\left((x+2)^2-10\right)^2}dx}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{(x+2)-1}{\left((x+2)^2-10\right)^2}dx}\) \(u=x+2\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{u-1}{\left(u^2-10\right)^2}du}\)

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{u-1}{\left(u^2-10\right)^2}du}\) \(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{u}{\left(u^2-10\right)^2}-\frac{1}{\left(u^2-10\right)^2}du}\) u sub and partial fractions for these two respectively

OpenStudy (solomonzelman):

well, r or any other letter sub (not u again)

OpenStudy (solomonzelman):

sorry for overtyping.

OpenStudy (anonymous):

Okay..i will try..thanks:)

OpenStudy (solomonzelman):

yw